Solve by elimination:
2x+3y=21
5x-2y=-14

Answer by funmath(2933)

Solve by elimination:
E1: 2x+3y=21
E2: 5x-2y=-14
Multiply E1 by 2 and E2 by 3. Add the two equations and the y's are eliminated and you can solve for x.
E1: 2(2x+3y)=2(21)---> 4x+6y=42
E2: 3(5x-2y)=3(-14)--->15x-6y=-42
:
4x+6y=42
15x-6y=-42
__________
19x+0=0
19x=0
19x/19=0/19
x=0
Substitute 0 in for x in either equation and solve for y. I'm choosing E2, Why don't you use E1 for fun. You'll find you get the same answer.
5(0)-2y=-14
0-2y=-14
-2y/-2=-14/-2
y=7
The solution is (x,y)=(0,7)
Happy Calculating!!!